OpenStudy (anonymous):

How do you know what to substitute in Solving Systems by Substitution? y=3x+2 7x-4y=7 how and why would you substitute?

OpenStudy (anonymous):

If you have a single variable in any equation, you can find an actual value for that variable. Substituting in a small equation into another equation is a method to find a value. So you've got this equation: y=3x+2. We know that's what y equals. All we have to do is take a look at the second equation and think about our next step. Well...We could just drag the little equation, y=3x+2, into 7x-4y=7, and see what we get! 7x - 4(3x+2) = 7 ...When (3x+2)=y See what happened? When I substituted in y=3x+2, I ended up with an equation with one single variable: X. Now, just solve! 7x - 12x - 8 = 7 - We used distribution here. -5x - 8 = 7 - Got rid of two like-terms. -5x = 15 - Added .-8 and 7 x = -3! - We found a value!!!! Now, use this information to find y! y = 3x + 2 y = 3(-3) + 2 y = -7 And we've got our answers. ^_^

OpenStudy (anonymous):

Always substitute the simplest looking equation into another given equation. I picked y=3x+2 instead of 7x-4y=7 because I knew right off the bat would y equaled.